In 1970, the consumer price index (CPI) for all items was 116.3. In 1990, it was 318.5. Find each of the following:
[
A. The cost of an item in 1990 to the nearest dollar that sold for $200 in 1970
B. The cost of an item in 1970 to the nearest dollar that sold for $200 in 1990
C. The percent change in the all items component between 1970 and 1990.
please help me im confused .

Answer by edjones(7311)

C) 318.5*100/116.3=273.86% change in 20 years.
.
A) 318.5/116.3*200=$547.72
.
B) Let x be the price in 1970.
318.5x/116.3=200
318.5x=200*116.3
318.5x=23260
x=23260/318.5
x=$73.03
.
Ed